OpenStudy (anonymous):

What is the sum of the arithmetic sequence 3, 11, 19 …, if there are 26 terms? 1,222 1,261 2,486 2,678

OpenStudy (anonymous):

d =8,a26=?, a26 =3+(25)8=2003 s=(26(3+203))/2=2678,answer

OpenStudy (anonymous):

2,678